OpenStudy (anonymous):

$850 at 8% for 8 years, compounded annually.

OpenStudy (jjuden):

first you find out what the interest is

OpenStudy (anonymous):

850(0.08)^8

OpenStudy (jjuden):

850*.08=68

OpenStudy (jjuden):

68*8=544 interest

OpenStudy (anonymous):

the options are 523.29 723.29 623.29

OpenStudy (anonymous):

thats why I'm confused because i got the same answer

OpenStudy (anonymous):

M=$(1+%)^time M=850(1+0.08)^8 M=850(1.08)^8 M=850x9.44 M=8024